OpenStudy (anonymous):

Expand the series and evaluate: Enter your answer as the following example: -1-2-3-4-5-6=-21 sum_(k=5)^10(-5k)

OpenStudy (amistre64):

whats your best effort give us?

OpenStudy (amistre64):

seeing this done a few times should give you an idea that we can build upon

OpenStudy (anonymous):

Ok @amistre64, please leave me alone

OpenStudy (amistre64):

ok, but good luck with it all :)

OpenStudy (anonymous):

Can someone help me please?

OpenStudy (amistre64):

im willing to help, but ....

OpenStudy (anonymous):

But what?

OpenStudy (amistre64):

you said you dont want me to help you

OpenStudy (anonymous):

I want help I just don't want to be criticized

OpenStudy (amistre64):

noones criticizing you. I just need to know how much you know on this subject.

OpenStudy (amistre64):

does the notation make sense for starters.

OpenStudy (amistre64):

\[\sum_{k=5}^{10}\]is a good place to start

OpenStudy (anonymous):

I know nothing

OpenStudy (amistre64):

then this might take a bit. It is simple enough tho. the symbol: \(\sum\) is a greek letter for "S" and indicated that we are going to be adding up stuff

OpenStudy (amistre64):

the number on the bottom tells us where we start at; and the number on top tells us where we end at: in this case we start with k=5 and end with k=10

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

i like to write it out to begin with like this: k : rule ------- 5 : -5k = -5(5) = -25 6 : -5k = -5(6) = -30 7 : -5k = -5(7) = -35 8 : -5k = -5(8) = -40 9 : -5k = -5(9) = -45 10: -5k = -5(10)= -50 does this make sense so far?

OpenStudy (amistre64):

\[\sum_{First}^{Last}(rule)\] using the rule, and stepping thru the numbers from first to last, we can generate all the numbers that are needed to be added up.